# TuneNSearch: a hybrid transfer learning and local search approach for solving vehicle routing problems

### v2

@@ 5.6 Analysis of the fine-tuning phase
For instances with 20 customers, POMO slightly outperforms TuneNSearch towards the end of training, however, for problems with 50 and 100 customers, POMO is inferior to TuneNSearch. Notably, the efficiency of our method is particularly evident at the beginning of training, since it benefits from prior knowledge gained during the pre-training phase. Furthermore, the performance gap between TuneNSearch and POMO appears to widen as problem sizes increase.
@@
Moreover, Table 11 presents the average training times for all models considered in this study. We note that the results of POMO and TuneNSearch are aggregated across all VRP variants. As shown, TuneNSearch requires about one-third of the computational time needed to train POMO from the beginning for each variant. It is important to highlight that the pre-trained MDVRP model incurs higher training time than POMO due to its use of a hidden dimension of $h_{x}$ = 256 instead of 128. Nonetheless, the strength of TuneNSearch lies in its capability to solve multiple VRP variants with the same pre-trained model and a fast fine-tuning phase.
@@ 5.7 Computational runtime complexity
Next, we investigate the computational runtime complexity of TuneNSearch on a set of benchmark instances, which are significantly large — up to 1000 nodes. The primary goal of this analysis is to understand how the integration of the local search algorithm scales with increasing problem sizes.
@@ 6 Conclusion
In this paper, we introduce TuneNSearch, a novel transfer learning method designed for adaptation to various VRP variants through an efficient fine-tuning phase. Our approach builds on the model architecture of Kwon et al. (2021) by enhancing the encoder with a residual E-GAT. While Lei et al. (2022) previously integrated this technique into the attention model (Kool et al., 2019), we extend its application to POMO, demonstrating that this extension improves the model’s ability to encode the problem’s features more effectively. Then, we pre-trained our model using MDVRP data, exploiting its complex graph-structured features. This strategy allows for a broader generalization across a variety of VRP variants, including both single- and multi-depot tasks. To evaluate the effectiveness of our approach, we compared it to an identical model pre-trained on the CVRP. The results demonstrate that while both methods perform similarly on single-depot variants, TuneNSearch substantially outperforms the CVRP-pre-trained model on multi-depot variants. To validate the learning process on the MDVRP, we compared TuneNSearch with MD-MTA, demonstrating superior results across all problem sizes while requiring much less training time. Finally, we integrated an efficient local search method to refine the quality of solutions generated by our model, leading to significant performance improvements with a small computational overhead. We also conducted an ablation study to assess the impact of varying the number of local search iterations. Our findings indicate that more constrained tasks tend to benefit more from this procedure.
To evaluate the generalization of TuneNSearch, we performed extensive experiments on numerous VRP variants and baselines. Experimental results on randomly generated instances show that TuneNSearch outperforms POMO, which is specialized for each variant, while requiring only one-fifth of the total training epochs. Moreover, TuneNSearch achieves performance comparable to OR-Tools guided local search procedure, often even outperforming it on many occasions, delivering results at a fraction of the computational time. We also provide results on benchmark instances of different VRP variants, where TuneNSearch outperforms other state-of-the-art neural-based models on most problems, narrowing the existing solution gaps. These findings demonstrate not only TuneNSearch’s strong cross-task generalization, but also its cross-size and cross-distribution generalization.
In addition to these results, we believe TuneNSearch holds significant potential for solving real-world problems. Many industries frequently face large, complex problems that demand rapid and near-optimal solutions. Traditional optimization algorithms struggle to generate high-quality solutions within polynomial time due to the NP-hard nature of routing problems, while conventional heuristics often result in large integrality gaps. TuneNSearch addresses these challenges by delivering good performance across various VRP variants while maintaining computational efficiency.
@@ 6.1 Limitations and future work
While this study offers insights into the development of generalizable neural-based methods, there are certain limitations to our approach. First, TuneNSearch is specifically designed to solve traditional VRPs, with the constraints outlined in Section 3.3 (or any combination of such constraints). Although it is possible to extend our method to other problems, such as the orienteering problem or prize collecting TSP, doing so would require manual adjustments to the Transformer architecture to accommodate new constraints. Second, TuneNSearch assumes that the objective function — minimizing the total distance traveled — remains unchanged across all scenarios. Modifying the objective during the fine-tuning phase may impact the learning process, as the neurons of the pre-trained model are already optimized for distance minimization. As such, in these scenarios, it is likely that TuneNSearch would require a higher number of fine-tuning epochs.
Looking ahead, we aim to extend TuneNSearch beyond traditional VRP variants to address a broader range of combinatorial problems. In particular, we plan to adapt our method to scheduling problems, which can be framed as variations of the TSP. Another promising direction is to further enhance the performance of TuneNSearch to reduce the integrality gaps even further. One possible approach could involve clustering training instances based on their underlying distributions and training specialized local models for each subset, which could potentially improve generalization. Alternatively, integrating decomposition approaches with learning-based methods to solve the VRP in a “divide-and-conquer” manner is another promising research direction.
@@ Appendix A Hyper-parameters sensitivity
First, we assessed the model’s sensitivity to the number of encoder layers by testing values of $L={3,4,5,6}$. Fig. 1 presents the average cost during pre-training for each value over 100 episodes. Interestingly, the model’s performance degraded when using 6 encoder layers, yielding results comparable to those obtained with 3 layers. In contrast, models with 4 and 5 encoder layers performed better, with the 5-layer configuration showing a slight edge over the 4-layer model.
@@ Appendix C Generalization on INCOM2024 benchmark instances
Instance Depots Customers PyVRP Ours MD-MTA POMO
Obj. Gap Obj. Gap Obj. Gap

### v3

@@ 5.6 Analysis of the fine-tuning phase
For instances with 20 customers, POMO slightly outperforms TuneNSearch towards the end of training, however, for problems with 50 and 100 customers, POMO is inferior to TuneNSearch. Notably, the efficiency of our method is particularly evident at the beginning of training, since it benefits from prior knowledge gained during the pre-training phase. Furthermore, the performance difference between TuneNSearch and POMO appears to widen as problem sizes increase.
@@
Moreover, Table 11 presents the average training times for all models considered in this study. We note that the results of POMO and TuneNSearch are aggregated across all VRP variants. As shown, TuneNSearch requires about one-third of the computational time needed to train POMO from the beginning for each variant. Nonetheless, the strength of TuneNSearch lies in its capability to solve multiple VRP variants with the same pre-trained model and a fast fine-tuning phase.
@@ 5.7 Computational runtime complexity
Next, we investigate the computational runtime complexity of TuneNSearch on a set of benchmark instances, which are significantly large, containing up to 1000 nodes. The primary goal of this analysis is to understand how the integration of the local search algorithm scales with increasing problem sizes.
@@ 6 Conclusion
In this paper, we introduce TuneNSearch, a novel transfer learning method designed for adaptation to various VRP variants through an efficient fine-tuning phase. Our approach builds on the model architecture of Kwon et al. (2021) by enhancing the encoder with an edge-aware mechanism. While Lei et al. (2022) previously integrated this technique into the attention model (Kool et al., 2019), we extend its application to POMO, demonstrating that this extension improves the model’s ability to encode the problem’s features more effectively. We first pre-train our model using MDVRP data, exploiting its complex graph-structured features. This strategy allows for a broader generalization across a variety of VRP variants, including both single- and multi-depot tasks. To evaluate the effectiveness of our approach, we compared it to an identical model pre-trained on the CVRP. The results demonstrate that while both methods perform similarly on single-depot variants, TuneNSearch substantially outperforms the CVRP-based model on multi-depot variants. To validate the learning process on the MDVRP, we compared TuneNSearch with MD-MTA, demonstrating superior results across all problem sizes while requiring much less training time. Finally, we integrated an efficient local search method to refine the quality of solutions generated by our model, leading to significant performance improvements with a small computational overhead. We also conducted experiments to assess the impact of increasing the number of local search iterations. Our findings indicate that more constrained tasks tend to benefit more from this procedure.
To evaluate the generalization of TuneNSearch, we performed extensive experiments on numerous VRP variants and baselines. Experimental results on randomly generated instances show that TuneNSearch outperforms POMO, which is specialized for each variant, while requiring only one-fifth of the total training epochs. Moreover, TuneNSearch achieves performance comparable to OR-Tools guided local search procedure, often even outperforming it on many occasions, delivering results at a fraction of the computational time. We also provide results on benchmark instances of different VRP variants, where TuneNSearch outperforms other state-of-the-art neural-based models on most problems, consistently achieving higher-quality solutions. These findings demonstrate not only TuneNSearch’s strong cross-task generalization, but also its cross-size and cross-distribution generalization.
@@ 6.1 Limitations and future work
While this study offers insights into the development of generalizable neural-based methods, there are certain limitations to our approach. First, TuneNSearch is specifically designed to solve traditional VRPs, with the constraints outlined in Section 3.3 (or any combination of such constraints). Although it is possible to extend our method to other problems, such as the orienteering problem or prize collecting TSP, doing so would require manual adjustments to the Transformer architecture to accommodate new constraints. Second, TuneNSearch assumes that the objective function (minimizing the total distance traveled) remains unchanged across all scenarios. Modifying the objective during the fine-tuning phase may impact the learning process, as the neurons of the pre-trained model are already optimized for distance minimization.
Looking ahead, we aim to extend TuneNSearch beyond traditional VRP variants to address a broader range of combinatorial problems. In particular, we plan to adapt our method to scheduling problems, which can be framed as variations of the TSP. Another promising direction is to further enhance the performance of TuneNSearch to improve solution quality even further. One possible approach could involve clustering training instances based on their underlying distributions and training specialized local models for each subset, which could potentially improve generalization. Alternatively, integrating decomposition approaches with learning-based methods to solve the VRP in a “divide-and-conquer” manner is another promising research direction.
@@ Appendix A Hyper-parameters sensitivity
First, we assessed the model’s sensitivity to the number of encoder layers by testing values of L=3,4,5,6L={3,4,5,6}. Fig. 1 presents the average cost during pre-training for each value over 100 episodes. Interestingly, the model’s performance degraded when using 6 encoder layers, yielding results comparable to those obtained with 3 layers. In contrast, models with 4 and 5 encoder layers performed better, with the 5-layer configuration showing a slight edge over the 4-layer model.
@@ Appendix C Generalization on INCOM2024 benchmark instances
Instance Depots Customers HGS-PyVRP Ours MD-MTA POMO
Obj. RPD Obj. RPD Obj. RPD